(6x+2)^1/3 - 4 = 3
Answer by stanbon(75887)

(6x+2)^1/3 - 4 = 3
(6x+2)^(1/3) = 7
Cube both sides to get:
6x+2 = 343
6x = 341
x = 56 5/6
================
Cheers,
Stan H.